### v1

# EmotiCrafter: Text-to-Emotional-Image Generation based on Valence-Arousal Model

Abstract
Recent research shows that emotions can enhance users’ cognition and influence information communication. While research on visual emotion analysis is extensive, limited work has been done on helping users generate emotionally rich image content. Existing work on emotional image generation relies on discrete emotion categories, making it challenging to capture complex and subtle emotional nuances accurately. Additionally, these methods struggle to control the specific content of generated images based on text prompts. In this work, we introduce the new task of continuous emotional image content generation (C-EICG) and present EmotiCrafter, an emotional image generation model that generates images based on text prompts and Valence-Arousal values. Specifically, we propose a novel emotion-embedding mapping network that embeds Valence-Arousal values into textual features, enabling the capture of specific emotions in alignment with intended input prompts. Additionally, we introduce a loss function to enhance emotion expression. The experimental results show that our method effectively generates images representing specific emotions with the desired content and outperforms existing techniques.
Figure 1: Generated images of the EmotiCrafter. Given a prompt and Valence-Arousal (V-A) values, our method can generate an emotional image that reflects the input prompt and aligns with the specified emotion values. By adjusting the V-A values, our method can also generate images that evoke discrete emotions, such as sadness, relaxation, anger, and amusement.
Yi He and Shengqi Dang contributed equally to this work.Nan Cao is the corresponding author.
1 Introduction
Emotions are fundamental to human experiences and play a critical role in shaping how people perceive and interact with the world. Research has shown that emotions affect memory [24, 20, 30, 5, 47] and comprehension [21, 42, 40], which are crucial for effective communication. As a result, content creators increasingly recognize the importance of incorporating emotions to enhance audience engagement and comprehension.
While research on visual emotion analysis is extensive [52, 18, 61], there is limited work on generating emotionally rich image content. Some early studies explored emotion generation techniques within specific domains such as facial expressions [2, 48] or landscapes [29]. EmoGen [53] generates images based on a given emotion tag (e.g., happy or sad), which pioneered the domain-free emotional image generation (EICG) task. However, it has two critical limitations: (1) the image is generated from an emotion tag instead of a text prompt, making the generated content difficult to control; (2) While the discrete emotion tags used in EmoGen are easy to understand, psychologists have not achieved a consensus on the emotion categories [63]. The limited scope of discrete emotion tags falls short of capturing nuanced emotions. In contrast, the continuous emotion models represent emotions through a continuous Cartesian coordinate system, allowing for a more precise capture of emotional complexity and subtlety [2]. Therefore, developing a general text-to-image model that reflects continuous emotional variations is desired.
In this paper, we introduce the task of continuous emotional image content generation (C-EICG) and present EmotiCrafter, an emotional image generation model that generates images based on text prompts and V-A values (i.e., two real numbers $(v,a)\in[-3,+3]$) defined in the Valence-Arousal (V-A) emotion model [37], the most frequently used continuous emotional model in psychology. An emotion-embedding network is designed to embed V-A values into the text prompts of a stable diffusion XL model (SDXL) [32]. This network enables a nonlinear embedding of V-A values into the features of the input text prompt to capture complex emotional context for image generation. To train the model, a dataset with 39,843 annotated images is prepared. Given an image, a training sample consists of a pair of V-A values about the image labeled by human annotators and two prompts, i.e., neutral prompt and emotional prompt generated by GPT-4 based on the image. The neutral [emotion] prompt describes the essential content of the image, whereas the emotional prompt describes the emotional information, such as the color, light, and texture, about the image. Finally, to enhance the generation results, we introduce a loss function that normalizes the V-A distribution and scales the residuals between the features of the neutral and emotional prompts provided in each training sample. In summary, our contributions are as follows:
  • We introduce a new task, i.e., continuous emotional image content generation (C-EICG), and propose a novel emotion-embedding network based on the transformer architecture. The network embeds the continuous Valence-Arousal values into the textual features of an input prompt for emotional image generation.
  • We develop a novel loss function that helps to enhance the emotional resonance of generated images. The loss function modifies the target features to learn more obvious emotional changes and takes the V-A distribution into account to counter the unevenness of the data.
  • We have constructed an emotional image dataset for training the proposed emotion-embedding network. Each training sample consists of a pair of unified V-A values, a neutral prompt, and an emotional prompt that are associated with the same image in the dataset.
2 Related work
In this section, we present a review of the related work, specifically focusing on visual emotion analysis, image emotion transfer, and conditional image generation.
2.1 Visual Emotion Analysis
Visual emotion analysis refers to the computational recognition and interpretation of human emotion [60] in visual media, such as images or videos. It has been a prominent research area, with most efforts focusing on the classification of discrete emotions [55, 51, 44, 4, 49]. However, discrete emotion categories limit the ability to capture the nuanced emotions [10], leading to an increased focus on continuous emotion analysis in images [62].
Much of the continuous emotion analysis remains centered on facial expression analysis [39, 27, 57, 13, 15]. While effective, facial analysis can overlook crucial contextual information that influences emotional interpretation. This drives studies to focus on emotions in objects or individuals within their environment, rather than just faces[26, 26, 16, 25, 18, 18]. For example, Kosti et al. [16] combined person-specific characteristics with the context of the scene to predict continuous emotional dimensions such as arousal, valence, and dominance. Kragel et al. [18] proposed EmoNet to extract visual features such as facial expressions, body posture, and scene elements from images to predict valence and arousal values. Recently, Mertens et al.[26] tested multiple backbones (such as ResNet, CLIP, DINO, etc.) for prediction of valence and arousal, demonstrating strong performance across various architectures.
Previous studies have achieved high accuracy in predicting continuous emotions in images, demonstrating a correlation between continuous emotions and visual elements. Building on this insight, our work moves beyond prediction to explore how continuous emotions can be actively embedded within generated content. Specifically, we introduce the task of continuous emotional image content generation (C-EICG) and propose a novel emotion-embedding network.
2.2 Image Emotion Transfer
Image Emotion Transfer (IET) focuses on editing the content of the images to evoke different emotions [22, 64, 65, 31, 45]. For instance, Peng et al. [31] achieved emotion transfer by adjusting color tones and texture-related features. Zhu et al. [65] introduced a method that separates high-level emotion-relevant features (e.g., object shapes and scene layout) from low-level emotion-relevant features (e.g., brightness). By applying GANs, they transferred emotions between images while preserving their original structure. Building on these methods, Weng et al. [45] proposed the Affective Image Filter, which uses a multi-modal transformer to process both text and image inputs. With the emergence of text-to-image models, IET has expanded into new applications based on instructive commands. For example, EmoEdit [54] used GPT-4V to build emotion factor trees that map abstract emotions to specific visual elements and employed the InstructPix2Pix model to apply emotion-driven content and color adjustments to images.
Current methods primarily extract specific features, often focusing on certain visual elements or emotional cues, which can limit the depth of emotional expression. In contrast, our method broadly learns a variety of emotion-influencing features through the emotion-embedding network, mapping continuous emotional variations at the textual level. Additionally, our model can accept natural language prompts as input, allowing for a more intuitive and flexible way to generate emotionally expressive images.
2.3 Conditional Image Generation
Conditional image generation aims to produce images that align with specific input conditions, such as text, depth maps, or sketches. Early explorations in this field utilized Variational Autoencoders (VAEs) [14] and Generative Adversarial Networks (GANs) [9]. Currently, textual conditional image generation models based on denoising diffusion [12], such as Stable Diffusion [35, 32] and DALL-E [34, 3], can generate exceptionally high-quality images. Besides, several studies have explored the use of other pieces of information as conditions for generation. The IP-adapter [56] integrates features from reference images through cross-attention, generating images consistent with attributes or identity. Techniques like Textual Inversion [8] and DreamBooth [36] can generate images in specific styles or domains by learning a concept. Models like ControlNet [58] and T2I-Adaptor [28] facilitate flexible conditional control using sketches, depth maps, and others.
However, there is limited research on the use of emotion as a condition. EmoGen [53] pioneered the generation of emotional image content (EIGC) by mapping emotional features to semantic features to create emotional images. Despite this, EmoGen struggles to understand natural language, limiting its capacity to effectively control specific content. Its reliance on discrete emotions also limits its practical applicability.
Our method bridges this gap by embedding continuous emotions into textual features, enabling image generation models to use both emotion and input prompts as conditions for creating emotional images.
Figure 2: Overview of our method. Specifically, we take the following steps. (a) We collect an image dataset annotated with V-A values, neutral prompts, and emotional prompts. These prompts are then encoded into features by prompt encoder $\mathcal{E}$. (b) Next, we design (b.1) an emotion-embedding network $\mathcal{M}$ to embed V/A values into textual features based on the transformer architecture, and (b.2) a specialized loss function to enhance the emotional resonance of generated images. The output of the mapping network serves as the condition for the image generation model $\mathcal{G}$ to generate emotional images.
Figure 3: Structure of Emotion Injection Block. It accepts hidden state $h_{i-1}$ as input and produces $h_{i}$ as output. The V-feature $e_{v}$ and A-feature $e_{a}$ represent the emotional features, which are injected through the cross-attention module.
3 Method
In this section, we introduce the technique details of EmotiCrafter. We begin with an overview of the proposed method, followed by an explanation of the network structure and key modules. Subsequently, we outline the training strategy. Finally, we describe the process of collecting and annotation of our dataset for network training.
3.1 Overview
As illustrated in Figure 2, our method generates emotional images $I_{emo}$ from two inputs: text prompts describing the desired image content and a pair of V-A values $(v,a)$ specifying the emotions. The process begins by encoding the input prompts into the prompt features $f_{n}$ using a prompt encoder $\mathcal{E}$. This features then pass through an emotion-embedding network $\mathcal{M}$ (Figure 2 (b.1)) to produce the emotional prompt features $\hat{f}_{e}$, embedded with emotion values $(v,a)$:
$\hat{f}_{e}$ $=\mathcal{M}(f_{n}|(v,a)).$ (1)
Finally, the emotional prompt features are input via the cross-attention mechanism in the SDXL $\mathcal{G}$ to generate the emotional images $I_{emo}=\mathcal{G}(\hat{f}_{e})$. To enhance the emotional expressiveness of generated images, we develop a loss function (Figure 2 (b.2)) that optimizes the network training by considering the V-A distribution of the training data and enhancing the residuals between the features of the emotional prompts and neutral prompts, compelling that generated images convey the desired emotions and content. To facilitate our emotion-embedding network, we have compiled an image dataset (Figure 2 (a)). This dataset pairs emotional images with both neutral prompts and emotional prompts, along with corresponding V-A values.
3.2 Emotion-Embedding Network
The emotion-embedding network $\mathcal{M}$ takes a pair of V-A values and a neutral prompt feature as input and outputs the corresponding emotional prompt feature (Figure 2 (b.1)). In particular, it first employs a V-A encoder to convert the V-A values into feature vectors. After that, an emotion injection transformer network (structure modified from GPT-2 [33] in our implementation) is used to infuse these features into the neutral prompt feature. Here, the transformer helps with capturing and preserving the textual context of the neutral prompt while injecting emotional features.
V-A Encoder. The V-A Encoder transforms a pair of V-A values into emotional features using two separate multilayer perceptrons (MLPs). The output V-features $e_{v}$ and the A-features $e_{a}$ are then transmitted into the emotion infusion process of the emotion injection transformer network.
Emotion Injection Transformer (EIT). The EIT is designed to effectively embed V/A-features into textual features through a modified GPT-2 architecture. The computing process of EIT consists of three steps: input projection, emotion injection, and output projection. First, we project the input neutral prompt feature $f_{n}$ into the transformer’s feature space:
$h_{0}=\text{P}_{\text{in}}(f_{n})+\text{PE},$ (2)
where $h_{0}$ represents the initial hidden state, $\text{P}_{\text{in}}(\cdot)$ is a linear projection layer, and PE denotes positional embedding [41] The core of EIT consists of 12 sequential Emotion Injection Blocks (EIBs) corresponding to 12 transformer blocks for emotion injection. Each block outputs a hidden state:
$h_{i}=\text{EIB}(h_{i-1},e_{v},e_{a}),\quad i\in\{1,\dots,12\}$ (3)
where $h_{i}$ is the output of the $i$-th EIB()($\cdot$)( ⋅ ). As shown in Figure 3, each EIB enhances the traditional transformer block through a cross-attention mechanism:
$h^{\prime}_{i}=\text{self-attn}(\text{LN}(h_{i-1}))+h_{i-1},$ (4)
$h^{(v)}_{i}=\text{cross-attn}(\text{LN}(h^{\prime}_{i}),e_{v})+h^{\prime}_{i},$ (5)
$h^{(v,a)}_{i}=\text{cross-attn}(\text{LN}(h^{(v)}_{i}),e_{a})+h^{(v)}_{i},$ (6)
$h_{i}=\text{fnn}(\text{LN}(h^{(v,a)}_{i}))+h^{(v,a)}_{i}),$ (7)
[…]
$\hat{f}_{r}=\text{P}_{\text{out}}(\text{LN}(h_{12})),$ (8)
The final emotional prompt feature is obtained by adding this residual to the original neutral prompt feature:
$\hat{f}_{e}=\hat{f}_{r}+f_{n}.$ (9)
3.3 Loss Function
The above emotion embedding network is trained by minimizing the averaged expectation of the difference between the predicted emotional prompt feature $\hat{f}_{e}=\mathcal{M}(f_{n}|(v,a))$ and the scaled target emotional prompt feature $f^{t}_{e}$, using the loss function described in Equation 10.
$\mathcal{L}=\frac{1}{n}\mathbb{E}\left(\frac{1}{d(v,a)}\|\hat{f}_{e}-f^{t}_{e}\|^{2}\right)$ (10)
where $n$ is the number of feature elements; $\mathbb{E}(\cdot)$ denotes the expectation; $d(v,a)$ is a density function that describes the distribution of V-A values in the training sample.
To effectively embed emotions and address the challenges posed by the uneven distribution of V-A values in the dataset, this loss function incorporates two key strategies to improve the model’s performance:
Scaled Residual Learning. To better capture pronounced emotional changes in generated images, we enlarge the target residuals:
$f^{t}_{e}=f_{n}+\alpha\underbrace{(f_{e}-f_{n})}_{\text{residual feature}},$ (11)
where $f_{e}$ is the emotional prompt feature, $f_{n}$ is the neutral prompt feature, and $\alpha$ is a scale factor, we set its value to 1.5 based on the ablation study.
V-A Density Weighting. To mitigate the effects of the imbalanced distribution of the training samples, we weigh the loss inversely proportional to the density of training samples in the V-A space. The density is estimated using Kernel Density Estimation (KDE) [6] with a Gaussian kernel, denoted as $d(v,a)$:
$d(v,a)=\frac{1}{n}\sum_{i=1}^{n}K_{H}((v,a)-(v_{i},a_{i})),$ (12)
where $K_{H}$ is a 2D Gaussian kernel with bandwidth $H$; $n$ is the number of training samples; $(v_{i},a_{i})$ are the V-A values of the $i$-th training sample. The bandwidth $H$ is selected using Silverman’s rule of thumb to provide optimal smoothing of the density estimation.
3.4 Dataset and Training
To train the emotion embedding network, an image caption dataset built on a set of annotated images is prepared. As shown in Figure 2, given an image, a sample in our dataset is a 3-tuple that consists of (1) a pair of V-A values about the image labeled by human annotators and two prompts, i.e., (2) neutral prompt and (3) emotional prompt. Both prompts are automatically generated using GPT-4, taking the image as the input. In particular, the neutral [emotion] prompt describes the essential content of the image, whereas the emotional prompt describes the emotional information, such as the color, light, and texture, about the image.
In particular, we collected OASIS [19], EMOTIC [17], and FindingEmo [26] to gather enough images with human-annotated V-A values. These images offer a diverse range of scenes beyond just facial images. After cleaning and merging, we compiled a final dataset of 39,843 images with V-A annotations. We applied linear scaling to standardize the varying V-A scales across these datasets, adjusting the values to a uniform range of $[-3,3]$ [26, 50]. Next, we applied prompt engineering to annotate the processed datasets with both neutral and emotional prompts, leveraging GPT-4’s vision capabilities to recognize essential content and express emotional nuances [7]. Please refer to the supplementary materials for further details.
The proposed emotion embedding network is trained based on two NVIDIA A800 GPUs using the above dataset. We utilize the AdamW [23] optimizer with a weight decay of 1e-5 and a learning rate of 1e-3. The training lasted for approximately 7 to 8 hours over a total of 200 epochs and 768 batch-size.
4 Experiments
In this section, we present both qualitative and quantitative comparisons, along with a user study and an ablation study, to evaluate the effectiveness of the proposed method.
Figure 4: Qualitative comparisons with baselines. These images are generated at varying V-A values, specifically -1.5, 0, and 1.5. Only our approach and the GPT-4+SDXL successfully generate images that clearly reflect emotional variations. Notably, our results show enhanced continuity, indicating superior controllability over continuous V-A values compared to the GPT-4+SDXL.
Figure 5: A more comprehensive comparison with the GPT-4+SDXL for the prompt “A group of people is skiing in a snow hill.” Our approach maintains continuity even under extreme V/A conditions. Conversely, GPT-4+SDXL displays noticeable discontinuities (e.g., in the V=3 column).
4.1 Comparisons
To estimate the effectiveness of the proposed method, we built four baselines based on existing techniques.
Baselines. We built baselines for comparison via two strategies: (1) inject the emotion features directly into the image generation modules in the SDXL such as UNet; (2) use the emotion features to change the input text prompt that controls the content of the generated image (similar to the proposed method).
As a result, four different baselines were built: (1) Cross Attention: inject the emotion features ($e_{v},e_{a}$) into the UNet in SDXL via its cross-attention mechanism based on IP-Adapter [56]; (2) Time Embedding: directly add emotion features ($e_{v},e_{a}$) to the time embedding of the UNet in SDXL. (3) Textual Inversion: use the text inversion technique [8] to embed emotion features ($e_{v},e_{a}$) into prompt templates with predefined emotion placeholders. (4) GPT-4+SDXL (GPT-SD): use GPT-4 [1] to rewrite the input text according to (v,a) values to generate an emotional SDXL prompt for image generation.
Qualitative Comparison. We evaluate the generated images based on three criteria: (1) the effectiveness of emotion embedding, (2) image-prompt similarity, and (3) the continuity of the image varies as V/A values change. Figure 4 illustrates that the baseline methods, Cross Attention, Time Embedding, and Textual Inversion, fail to reveal the continuous change of emotions. Notably, Textual Inversion tends to generate images with a persistent purple tint, which significantly limits its emotional expressiveness. Only our method and the GPT-4+SDXL successfully generate images that clearly reflect emotional variations. It also demonstrates that all methods generate content faithfully to the given prompts. For image continuity, we focus on comparing our method and GPT-4+SDXL, as they are the only two that successfully generate emotional variations. Our method shows enhanced continuity, indicating superior controllability over continuous V-A values. Figure 5 offers a more comprehensive comparison, revealing that our approach maintains continuity even under extreme V/A conditions. Conversely, GPT-4+SDXL displays noticeable discontinuities (e.g., in the V=3 column).
Quantitative Comparison. We compare our method against several baselines using the following metrics: (1) V/A-Error evaluates the absolute error between the predicted V/A values of the generated images and the input V/A values. (2) CLIPScore [11] assesses the similarity between the input text and the generated images. (3) CLIP-IQA [43] leverages a pre-trained CLIP model to evaluate image quality without requiring reference images. (4) LPIPS-Continuous utilizes the Learned Perceptual Image Patch Similarity [59] to measure the continuity of the change in image as V / A changes.
A-Error $\downarrow$ V-Error $\downarrow$ CLIPScore $\uparrow$ CLIP-IQA $\uparrow$
Cross Attention 1.923$\pm$1.153 2.080$\pm$1.438 26.266$\pm$2.381 0.949$\pm$0.046
Time Embedding 1.941$\pm$1.168 2.031$\pm$1.348 26.566$\pm$2.125 0.786$\pm$0.164
Textual Inversion 1.958$\pm$1.188 1.923$\pm$1.170 22.346$\pm$3.594 0.370$\pm$0.111
GPT-4+SDXL 1.860$\pm$1.090 1.517$\pm$1.060 25.907$\pm$1.949 0.906$\pm$0.066
Ours 1.828$\pm$1.085 1.510$\pm$1.074 23.067$\pm$2.655 0.881$\pm$0.099
Table 1: Quantitative results comparing emotion accuracy, prompt fidelity, and image quality across different baselines (132 prompts × 5 V/A values × 25 images = 3,300 images per method). Our method achieves the highest performance in emotion accuracy, while maintaining comparable results in prompt fidelity and image quality, with a minor decrease due to the fact that altering emotional content will inherently affect semantic alignment.
As shown in Table 1, our method achieves the best (lowest) V/A-Error on average. Cross Attention and Time Embedding achieve the best (highest) CLIP-IQA and CLIPScore respectively. However, these methods fail to generate emotional images, as shown in Figure 4. Our method exhibits a slight decrease in CLIPScore relative to the baselines. We believe this is due to the fact that altering emotional content will inherently affect semantic alignment. At the same time, CLIP-IQA value shows that our method could generate high-quality images. Our method demonstrates superior continuity when compared to that of GPT-4 + SDXL (Table 2).
Ours GPT-4+SDXL
[…]
Table 2: The continuity comparison between our method and GPT-4+SDXL (baseline).
4.2 User Study
We conducted two user studies with 20 college students to evaluate the effectiveness of our method by comparing it to GPT-4+SDXL (the baseline).
Ours GPT-4+SDXL
A-Ranking Consistency $\uparrow$ 0.759±0.273 0.165±0.379
V-Ranking Consistency $\uparrow$ 0.887±0.245 0.584±0.259
A-Error $\downarrow$ 1.327±1.120 2.029±1.446
V-Error $\downarrow$ 0.692±0.682 1.229±1.026
Emotion Consistency $\uparrow$ 4.215±0.715 3.525±1.065
Emotion Smoothness $\uparrow$ 4.240±0.828 3.195±1.163
Table 3: User study. Our method outperformed the baseline in all aspects.
Experiment I. The first experiment tested whether the emotion represented by the generated images aligned with a human’s perception. To this end, two collections of images were prepared respectively for testing the perception of A and V values. Each collection consists of 20 sets of images with 10 sets generated by our method and the other 10 sets generated by the baseline. Each set consisted of 5 images with different A or V values whose orders were randomized. The participants were asked to reorder these 5 images respectively based on the perceived intensity of V or A. Kendall’s $\tau_{b}\in[-1,+1]$ was computed to assess the alignment between the order given by participants and the ground truth order, where $\tau_{b}=1$ indicated a full match of the orders. During the experiment, We also asked participants to use their perceived emotions to estimate and label the V and A values of each image. The absolute error between the results and the ground truth was calculated.
Experiment II. The second experiment was designed to test whether the generated images were able to reveal the contentious change of emotions (i.e., V-A values). To this end, we generated 20 sets of images (10 based on our method and the other 10 based on baseline). As shown in Figure 5, each set consists of 25 images generated based on V-A values gradually changed from -3 to +3. Participants were asked to rate each image set (using a 5-point Likert scale) regarding two aspects: (1) the consistency between the change of V-A values and the change of image content; (2) the smoothness of the change of the image content.
Analysis & Results. We conducted a Shapiro-Wilk test [38] to assess normality and applied the Wilcoxon Signed Rank test [46] to evaluate the significance of all the results. The results showed that our method outperformed the baseline in all aspects (Table 3) and showed a statistical significance regarding V/A Ranking, V/A Error, Consistency, and Smoothness (Figure 6).
Figure 6: Significance test. (a) V/A Ranking: Kendall’s $\tau$b, (b) V/A Error: absolute error, (c) Emotion consistency, and Emotion smoothness. Significance levels: *p<0.05, **p<0.01, ***p<0.001
Figure 7: Abalation Study. Images are generated from the prompt “An oil painting shows an astronaut.” As $\alpha$ increases, image-prompt similarity decreases, while emotional variations increase. The usage of $d(a,v)$ enhances the accuracy of emotional changes in the generated images.
4.3 Ablation Study
We conducted ablation studies to estimate the effectiveness of the proposed loss function.
Figure 8: The effectiveness of the scaling factor $\alpha$. The method with $\alpha=1.5$ ($\bigstar$) surpasses the performance indicated by the regression line (- - -).
Effectiveness of the Scaling Factor $\alpha$. We tested the changes of CLIPScore and V/A-Error while the scaling factor $\alpha$ varies. The results (Figure 8) showed that when $\alpha$ increased, both the CLIPScore and V/A-Error decreased (CLIPScore getting worse, V/A-Error getting better). This finding can also be observed from the examples shown in Figure 7. We set $\alpha=1.5$ as a trade-offer regarding the experiment results. However, users still can choose different $\alpha$ values to customize their output.
A-Error $\downarrow$ V-Error $\downarrow$ CLIPScore $\uparrow$
Ours 1.828±1.085 1.510±1.074 23.067$\pm$2.655
w/o $d(v,a)$ 1.829±1.083 1.546±1.082 21.977$\pm$0.066
Table 4: The effectiveness of $d(v,a)$.
Effectiveness of the Density Weighting $d(v,a)$. We compare our method with a variant that removed $d(v,a)$ from the training process. As shown in Table 4, incorporating $d(v,a)$ into the loss function positively affects both the CLIPScore and V/A-Error, which can also be easily found in Figure 7.
4.4 Discussion
Although our method achieved promising results, it still has some limitations that are worth mentioning: First, we observe that controlling image generation based on arousal is more challenging compared to valence. This aligns with prior research in visual emotion analysis, which has similarly reported that arousal is harder to predict, largely due to lower agreement among annotators when assessing arousal levels [26]; Second, our approach often generates images featuring human activities, even when such activities are not specified in the prompts. This tendency is likely attributable to the limited representation of non-human scenes in the training data and could be addressed by incorporating a greater variety of non-human scenarios into the dataset; Third, our method occasionally adjusts users’ input prompts to better align with the specified emotional prompts, resulting a slight shift of the prompt meaning, thus effects the image generation. We believe this issue could be addressed by adding a term for semantic preserving in the loss function.
5 Conclusion
In this paper, we introduce a new task, i.e., continuous emotional image content generation (C-EICG), and present EmotiCrafter, a novel method for generating emotionally expressive images based on continuous emotional values (Valence-Arousal). Specifically, we introduce an emotion-embedding network that integrates V-A values into textual features. Extensive experiments demonstrate the effectiveness of our approach, showing that it can reliably generate images that align with both user-provided prompts and specified emotions described by V-A values. We believe this work will catalyze further advancements in affective computing and image generation. To encourage future research, we will release our code and data, hoping it will inspire new developments in the field.